OpenStudy (anonymous):

(x-6)^2=81. can someone help me solve this please.

OpenStudy (anonymous):

x^2 - 12x + 36 = 81 x^2 - 12x = 45 Now you can solve it using formula for quadratic equation.

OpenStudy (anonymous):

this is what i have so far \[x-6=\sqrt{81} = x-6=9\]

OpenStudy (anonymous):

almost tena, just remember if you "force" a square root you need the ± of the number

OpenStudy (anonymous):

is it -3,-15

OpenStudy (anonymous):

x−6=9 x−6=-9 x = 15 x = -3

OpenStudy (anonymous):

see i cant find that symbol using the mathlab..oh wow look i was almost right!!! lol

OpenStudy (anonymous):

you have no clue how long i have been working on this math.. i dont understand NONE of it

OpenStudy (anonymous):

On a mac to get that symbol you hold "alt shift +" and it makes the ± On the other note you had most of the right procedure, the only thing you missed was that √81 = ±9

OpenStudy (anonymous):

oh im not on a mac but ill try and see what it does on mine. nope it dont work lol